OpenStudy (anonymous):

If the equation of an ellipse is in standard form and the denominators are equal, then the ellipse is a circle. Is this true or false?

OpenStudy (phi):

do you know the equation of a circle? Start with the equation of an ellipse: \[ \frac{(x-h)^2}{a^2} + \frac{(y-k)^2}{b^2 } = 1 \] make the denominators equal. OK, let b = a so we have \[ \frac{(x-h)^2}{a^2} + \frac{(y-k)^2}{a^2 } = 1 \] or \[ \frac{(x-h)^2+(y-k)^2}{a^2} =1 \] if you multiply both sides by a^2 you get \[ (x-h)^2+(y-k)^2 = a^2 \]

OpenStudy (anonymous):

Thank you!